OpenStudy (anonymous):

ordered pair to solve -x+4y=9 (-1,__)

OpenStudy (anonymous):

-1,2

OpenStudy (rosedewittbukater):

I'm guessing that means x=-1, so plug that in. -(-1) + 4y = 9. Then solve for y. 1 +4y = 9 4y = 9-1 4y = 8 y = 2 so (-1, 2)